# Explainability of Text Processing and Retrieval Methods: A Survey

### v2

@@ 4.1. Using probing classifiers to construct global explanations
While a number of studies (Köhn, 2015; Gupta et al., 2015; Hupkes et al., 2018) have used probing methods to evaluate hypotheses like “neural models learn syntactic structures”, Belinkov (2022) highlights the drawbacks of probing methods, some of which are described below.
@@
4.2. LIME: model-agnostic, local explanations
@@
LIME (Ribeiro et al., 2016) is one of the earliest methods to be proposed in the current wave of research on explainability. Given an input $I$, LIME generates a set $\pi_{S}$ of slightly perturbed instances of $I_{S}$ (recall that $I_{S}$ denotes some interpretable representation of $I$: for a document, $I_{S}$ may be a binary term-vector), by randomly setting a small fraction of the nonzero elements of $I_{S}$ to zero. For each perturbed instance, the corresponding complex feature vector $I_{C}$ is also constructed; together, these comprise $\pi_{C}$. A simple classifier, e.g., a linear model, is trained on $\pi_{S}$ to mimic (and thus “explain”) $\mathcal{M}_{C}$ on $\pi_{C}$. No assumptions are made about $\mathcal{M}_{C}$. LIME is thus a model-agnostic (§3.1) method that provides simple surrogates (§3.4.1) as local (§3.3) explanations.
4.3. Feature attribution as model-agnostic, local explanations
@@
As discussed in § 3.4.2, the objective of feature attribution is to quantify the contribution of a particular feature towards the output of $\mathcal{M}_{C}$ on $I$. Possibly the simplest feature attribution technique is occlusion (Zeiler and Fergus, 2014; Li et al., 2016) (sometimes referred to as the leave one out method). This estimates the importance of a group of features by setting all these features to zero, and measuring the resulting decrease in prediction accuracy. [^3] This general idea has been used in a wide variety of contexts where effects have to be measured. Since this approach only examines how the output of the model is affected by changes in the input, it is model-agnostic. Two other commonly-used families of feature attribution methods are discussed below. A library implementing many of these methods is available from https://captum.ai/, and may be used to interpret PyTorch models.
22todo: 2Possible to fit deeplift (deeplift) into one of the two following subsections?
4.3.1. Gradient based techniques
@@
Sundararajan et al. (2017) proposed Integrated Gradients, a well-known method in this class of approaches. To quantify the importance of each feature, $I_{C}$ (the feature vector of an input instance $I$) is compared to a “baseline” ($I^{(0)}$, say) that corresponds to a very simple input. For example, if the input is an image, the baseline could be a black image; when studying word embeddings, the zero vector would be a possible baseline. We consider $m$ additional instances by linearly interpolating between $I_{C}$ and $I^{(0)}$. If $\mathbf{i}_{j}$ denotes the $j$-th feature or dimension of $I_{C}$, then its importance (or effect on the prediction) is measured in terms of its integrated gradient $IG_{j}(I_{C},I^{(0)})$, which depends on two factors:  (i) the difference between $\mathbf{i}_{j}$ and its corresponding value $\mathbf{i}^{(0)}_{j}$ in the baseline $I^{(0)}$,  and  (ii) $\frac{\partial f(I_{C})}{\partial\mathbf{i}_{j}}$, the gradient of the prediction $f$ with respect to $\mathbf{i}_{j}$. Instead of using the value of the gradient at a single point, we use the average of the gradients computed at the $m$ interpolation points between $I_{C}$ and $I^{(0)}$. More concretely,
@@
4.3.2. Shapley
The Shapley value (Shapley, 2016) was originally proposed in 1952 in the context of Cooperative Game Theory to quantify the contribution of a player to the outcome of a game. When a team of $k$ players participates in a game and receives a payout, the objective of Shapley is to distribute the payout among team members in a fair way, i.e., in accordance with their respective contribution. Lundberg and Lee (2017) first proposed the use of Shapley values to explain ML models: each feature may be regarded as a player, the model itself is the game, the prediction of the model is the payout, and the Shapley value for each feature measures its contribution towards the overall prediction. The time complexity of computing Shapley values according to the original formulation is exponential, since it involves examining all possible subsets of features. Lundberg and Lee (2017) formulate approximate versions, viz., Kernel SHAP, Tree SHAP (tree based approximation of SHAP), and Deep SHAP, that may be computed in reasonable time.
@@ 5. Evaluating explanations
The subsequent sections describe diverse approaches that have been proposed to explain neural models for text representation/indexing and ranking. In order to measure advances in this area, however, evaluation is possibly the biggest issue that needs to be resolved. Humans constitute the primary target of explanations; the diversity of this target group makes it challenging to formulate an evaluation framework. The consumers of explanations range from lay users, through persons responsible for checking compliance with ethical/legal requirements and domain experts who may not be familiar with the principles underlying IR/NLP models, to IR/NLP researchers and practitioners, whose primary objective may be failure analysis, or to gain insights that guide further development. Of course, this diversity is also true of the users of IR/NLP systems, and the original Cranfield paradigm for evaluating IR had several drawbacks. Some of these have been addressed, while some remain, but over the years, a generally accepted protocol for evaluating IR/NLP systems has emerged. Importantly, a number of large test collections have been created for many different tasks, and their role in advancing the field cannot be overemphasised.
@@
Even though similar, widely accepted protocols and test collections for ExIR are yet to emerge, two broad evaluation paradigms may be identified. The first involves anecdotal evaluation by humans. This seems unavoidable, since most reearch groups do not have the resources to conduct carefully designed, large-scale user studies. While this sort of evaluation may be enough to establish the plausibility of explanation methods, it is not rigorous. The second approach to evaluation utilises different quantitative notions of similarity. When explanations are in the form of simple surrogates that mimic complex models (see §3.4), evaluation is typically based on measuring the similarity between the behaviours of the explanation and the complex model, using standard metrics like rank correlation. Likewise, feature attribution is generally evaluated by measuring the difference in the performance of a model when it is given the original inputs, and when it is given a simplified representation consisting of just the putatively important features. Rationale-based approaches may be evaluated similarly; in cases where human annotated rationales are available, explanations may also be evaluated by measuring the overlap between the generated rationales and the ground truth.
33todo: 3complete this section
@@ 6. Document Ranking
Neural models for document ranking are generally categorised as representation-based or interaction-based. Like traditional retrieval models, representation-based models also encode queries and documents as vectors; however, in contrast to the sparse vectors used by traditional approaches, these vectors are dense. On the other hand, interaction-based models capture complex interactions across query-term–document-term pairs. This section provides an overview of efforts to explain both categories of neural ranking methods, but first, we provide a brief review of some foundational attempts to understand traditional IR models (§ 6.1). These early studies were all targeted towards global explanations. Next, in Section 6.2, we discuss the Axiomatic Framework (Fang et al., 2004) — a cornerstone of a substantial body of work on explaining IR — and its recent applications. In Sections 6.36.5, we review other approaches to providing global explanations of Neural Ranking Models (NRMs). Finally, in Sections 6.66.8, we turn to studies devoted to local explanations. Tables 3 and 4 summarise the articles covered in these sections, according to whether they address global or local approaches. We are also preparing a Python package, ir_explain (Saha et al., 2024), that provides implementations of well-known post-hoc explanation strategies (including pointwise, pairwise, and listwise methods) within a single framework.
@@ 6.1. Understanding traditional models: the vector space model, and language modeling
Traditional IR systems typically represent documents using terms, i.e., single words (or word stems), and sometimes, additionally, groups of words that have special significance (these are often loosely referred to as phrases), along with their associated weights. The term weights (or feature weights) are calculated using either intuitive heuristics, or on the basis of various mathematical models. Traditional IR models may thus be regarded as explainable by design.
Nonetheless, some early work focused on developing a better understanding of these models. Singhal (1997) carefully examined traditional term-weighting formulae used in the Vector Space Model, explained why they do not work as well as more effective statistical models like BM25, proposed ways to modify the formulae in order to address these flaws, and demonstrated that the modified formulae do indeed yield significant improvements in retrieval effectiveness. Besides conventional text collections, Singhal also considered noisy text collections, such as those generated by running Optical Character Recognition (OCR) on scanned document images, and explained why it is inadvisable to use cosine normalisation for such collections. Kamps et al. (2005) studied similar issues within the context of retrieving XML elements from a collection of XML documents, to explain the importance of length normalisation in XML retrieval. On a related note, Smucker and Allan (2006) compared two well-known smoothing techniques used in Language Modeling, and presented a different explanation for why Dirichlet smoothing outperforms Jelinek-Mercer smoothing, compared to the perspective provided by Zhai and Lafferty (2004). In terms of the criteria introduced in § 3, the explanations provided in (Singhal, 1997; Kamps et al., 2005; Smucker and Allan, 2006) are all model-aware, global explanations. We conclude this section by mentioning I-REX (Roy et al., 2019), an interactive tool that is intended to facilitate further studies along these lines by providing a white-box view into traditional IR models. For example, given a retrieval model, and a query $Q$, the compare command in I-REX provides a pairwise explanation (§ 3.3.1) of the ranks of documents $D_{i},D_{j}$, by displaying the query terms present in $D_{i},D_{j}$ along with their term frequencies, collection frequencies, and overall contribution to the query-document scores.
@@ 6.2. The axiomatic framework: improving and explaining IR models
There are certain common principles underlying the different term-weighting formulae that are commonly used, even though they are derived using different mathematical models. In a series of studies (Fang et al., 2004; Fang and Zhai, 2006; Fang et al., 2011), Fang et al. consolidated these principles, and formalised them within an axiomatic framework, as a set of constraints, or axioms, that term-weighting schemes must satisfy, in order to be effective. This framework provides a natural basis for pairwise explanations (§3.3.1), i.e., for reasoning about why a particular document $D_{i}$ is ranked above $D_{j}$ by a ranking model for a given query $Q$. It has been used to explain why certain retrieval models work better than other models, as well as to formulate improved versions of existing retrieval models by better aligning them with IR axioms. (Clinchant and Gaussier, 2009b, a). Later efforts have focused on refining the framework by incorporating additional axioms (Tao and Zhai, 2007; Fang and Zhai, 2006), as well as expanding it to cover pseudo-relevance feedback methods (Clinchant and Gaussier, 2013). Table 2 lists some of the most popular axioms that have been mentioned in the literature. The page at https://www.eecis.udel.edu/~hfang/AX.html is an excellent resource on this topic.
@@
SemanticMatches STMC1 (Fang and Zhai, 2006) favour a document containing terms that are more semantically related to query terms
@@
STMC2 (Fang and Zhai, 2006) avoid over-favouring semantically similar terms
@@
STMC3 (Fang and Zhai, 2006) favour a document containing more terms that are semantically related to different query terms
@@
TSFC (Pang et al., 2017) a document having higher exact match should get higher score if their semantic similarity score matches
@@
Term Proximity TP (Tao and Zhai, 2007) favour a document in which query terms appear in closer proximity
@@
Table 2. Popular axioms proposed in the past literature on axiomatic IR research (§ 6.2), adapted from (Fang et al., 2011; Câmara and Hauff, 2020).
@@
The axiomatic framework has been applied in recent times to analyse and understand neural IR models in terms of the common principles underlying traditional IR models (which are, in a sense, explainable by design). Some of these efforts investigate the extent to which various IR models (neural, as well as traditional) satisfy axiomatic constraints. Others try to approximate the ranking generated by a retrieval model via an explainable surrogate, a la LIME (§4.2).
@@
Pang et al. (2017) investigated the difference between the features used in traditional models (BM25 and LMJM) and two pre-BERT NRMs: ARC-I (Hu et al., 2014), a representation based model, and MatchPyramid (Pang et al., 2016b), an interaction based model. Their experiments on the Robust04 and LETOR 4.0 (Qin et al., 2010) datasets suggest that features automatically learned by NRMs are more robust compared to the word-based features used by BM25 and LMJM, in the sense that the performance of NRMs is less affected than that of traditional models when some features are removed. They also confirmed that interaction-based models appear to focus on query related words, while representation-based models seem more concerned with general topic related words, such as those generated by a state-of-the-art LDA based model. Interestingly, Pang et al. noted that ARC-I and MatchPyramid performed poorly on LETOR 4.0 and even worse on Robust04, compared to BM25 and LMJM. Their failure analysis suggests that these NRMs do not satisfy the TDC axiom (Table 2). In addition, due to limitations on memory and time, documents were sometimes truncated to some fixed length, e.g., 500 words. This truncation may reduce the number of query terms present in a document, resulting in a violation of the TFC1 axiom. It was also suggested that semantic matching may be introducing too much noise into the scoring function. A new axiomatic constraint TSFC (Table 2) was proposed, which states that if two documents are of equal length, and have equivalent semantic term matching signals, the document having a larger exact match should be ranked higher. Furthermore, the authors argued that the size of Robust04 is not suitable for training NRMs. 44todo: 4how the words are generated in arc-i pooling e? second version e.
@@
Following the trend of creating diagnostic datasets for NLP and Computer Vision applications, Rennings et al. used selected axioms to construct a similar dataset for IR. Query-document-pair triplets that satisfy TFC1, TFC2, TDC (cf. Table 2) were sampled from the WikiPassageQA (Cohen et al., 2018b) corpus. Artificial documents were constructed to create triplets satisfying LNC2. This diagnostic dataset was used to evaluate the axiomatic efficacy (the fraction of axiomatic constraints that is satisfied by a retrieval model) of both traditional models (BM25 and LMDIR), and neural models like DRMM (Guo et al., 2016), aNMM (Yang et al., 2016), Duet (Mitra et al., 2017) and MatchPyramid (Pang et al., 2016b). The authors observed that statistical models outperform neural models in terms of axiomatic efficacy, but were significantly outperformed by DRMM and aNMM in terms of retrieval effectiveness. In later work, Câmara and Hauff (2020) extended this idea by considering five additional axioms (LNC1, STMC1, STMC2, STMC3, TP). Once again, empirical results showed that DistilBERT (Sanh et al., 2019) significantly outperforms traditional query likelihood methods, but it does not perform as well in terms of axiomatic efficacy. It is particularly counter-intuitive that this observation applies to semantic matching axioms, e.g., STMC1, as well. These results suggest that existing axioms, including the semantic matching axioms, are not adequate for analysing NRMs based on BERT and its variants, and need to be formulated more carefully.
@@
Völske et al. (2021) provided post-hoc explanations of IR systems by constructing an explanation model to approximate the ranking generated by a retrieval model. Recall that each axiom induces relative orderings for certain document pairs; e.g., axiom $A_{i}$ may suggest that $D_{j}$ should be ranked above $D_{k}$. The explanation model (a random forest) may be regarded as a weighted aggregation of these orderings. In the words of the authors, “(a) the explanation model’s parameters reveal the degree to which different axiomatic constraints are important to the retrieval model under consideration, and (b) the fidelity with which the initial ranking can be reconstructed can point to blind spots in the axiom set, which can help to uncover new ranking properties yet to be formalized.” Twenty axioms were used in this study. The behaviour of three statistical models (BM25, TF-IDF, PL2) and five neural models (MP-COS (Pang et al., 2016b), DRMM (Guo et al., 2016), PACRR-DRMM (McDonald et al., 2018), BERT-3S (Akkalyoncu Yilmaz et al., 2019), DAI-MAXP (Dai and Callan, 2019)) were studied using the MSMARCO (Nguyen et al., 2016) and Robust04 collections. [^4] Table 9 summarises information about some datasets commonly used in IR explainability studies. The authors observed that it is difficult to explain the relative ordering of two documents if they have similar retrieval scores, but a big difference in scores can be reliably explained in terms of the axioms. They also noted that explanation effectiveness is comparable for traditional and neural models (in this study, explanation efficacy is measured in terms of how accurately the explanation model can replicate a retrieval model’s ranking), but it was higher for the Robust04 collection, as compared to MSMARCO.
@@
The use of the axiomatic framework in these studies (Rennings et al., ; Câmara and Hauff, 2020; Völske et al., 2021) draws on earlier work by Hagen et al. (2016) that is not directly related to explainability. Hagen et al. were likely the first to formulate relaxed versions of axioms that permitted them to be used for practical analysis,. Axiom TFC1 (cf. Table 2), for example, induces an ordering on a document pair based on the term-frequencies of a query term in the two documents, provided the two documents are of equal length. To be applicable in practice, this equality condition was replaced by an approximate equality constraint: axiom TFC1 was applied to a document pair if the relative difference in the lengths of the two documents was at most 10%. The work discussed above (Rennings et al., ; Câmara and Hauff, 2020; Völske et al., 2021) uses such relaxed formulations of various axioms.

### v3

@@ 4.1. Using probing classifiers to construct global explanations
While a number of studies (kohn-2015-whats; gupta-etal-2015-distributional; hupkes-rnn) have used probing methods to evaluate hypotheses like “neural models learn syntactic structures”, 10.1162/coli_a_00422 highlights the drawbacks of probing methods, some of which are described below.
@@
4.2. Model-agnostic, local explanations
4.2.1. LIME
@@
LIME (lime) is one of the earliest methods to be proposed in the current wave of research on explainability. Given an input $I$, LIME generates a set $\pi_{S}$ of slightly perturbed instances of $I_{S}$ (recall that $I_{S}$ denotes some interpretable representation of $I$: for a document, $I_{S}$ may be a binary term-vector), by randomly setting a small fraction of the nonzero elements of $I_{S}$ to zero. For each perturbed instance, the corresponding complex feature vector $I_{C}$ is also constructed; together, these comprise $\pi_{C}$. A simple classifier, e.g., a linear model, is trained on $\pi_{S}$ to mimic (and thus “explain”) $\mathcal{M}_{C}$ on $\pi_{C}$. No assumptions are made about $\mathcal{M}_{C}$. LIME is thus a model-agnostic (Section 3.1) method that provides simple surrogates (Section 3.4.1) as local (Section 3.3) explanations.
4.2.2. Feature attribution
@@
As discussed in Section 3.4.2, the objective of feature attribution is to quantify the contribution of a particular feature towards the output of $\mathcal{M}_{C}$ on $I$. Possibly the simplest feature attribution technique is occlusion (10.1007/978-3-319-10590-1_53; DBLP:journals/corr/LiMJ16a) (sometimes referred to as the leave one out method). This estimates the importance of a group of features by setting all these features to zero, and measuring the resulting decrease in prediction accuracy. [^3] This general idea has been used in a wide variety of contexts where effects have to be measured. Since this approach only examines how the output of the model is affected by changes in the input, it is model-agnostic. Two other commonly-used families of feature attribution methods are discussed below. A library implementing many of these methods is available from https://captum.ai/, and may be used to interpret PyTorch models.
Gradient based techniques
@@
10.5555/3305890.3306024 proposed Integrated Gradients, a well-known method in this class of approaches. To quantify the importance of each feature, $I_{C}$ (the feature vector of an input instance $I$) is compared to a “baseline” ($I^{(0)}$, say) that corresponds to a very simple input. For example, if the input is an image, the baseline could be a black image; when studying word embeddings, the zero vector would be a possible baseline. We consider $m$ additional instances by linearly interpolating between $I_{C}$ and $I^{(0)}$. If $\mathbf{i}_{j}$ denotes the $j$-th feature or dimension of $I_{C}$, then its importance (or effect on the prediction) is measured in terms of its integrated gradient $IG_{j}(I_{C},I^{(0)})$, which depends on two factors:  (i) the difference between $\mathbf{i}_{j}$ and its corresponding value $\mathbf{i}^{(0)}_{j}$ in the baseline $I^{(0)}$,  and  (ii) $\frac{\partial f(I_{C})}{\partial\mathbf{i}_{j}}$, the gradient of the prediction $f$ with respect to $\mathbf{i}_{j}$. Instead of using the value of the gradient at a single point, we use the average of the gradients computed at the $m$ interpolation points between $I_{C}$ and $I^{(0)}$. More concretely,
@@
Shapley
The Shapley value (shapley) was originally proposed in 1952 in the context of Cooperative Game Theory to quantify the contribution of a player to the outcome of a game. When a team of $k$ players participates in a game and receives a payout, the objective of Shapley is to distribute the payout among team members in a fair way, i.e., in accordance with their respective contribution. deep-shap first proposed the use of Shapley values to explain ML models: each feature may be regarded as a player, the model itself is the game, the prediction of the model is the payout, and the Shapley value for each feature measures its contribution towards the overall prediction. The time complexity of computing Shapley values according to the original formulation is exponential, since it involves examining all possible subsets of features. deep-shap formulate approximate versions, viz., Kernel SHAP, Tree SHAP (tree based approximation of SHAP), and Deep SHAP, that may be computed in reasonable time.
@@ 5. Evaluating explanations
The following sections describe diverse approaches that have been proposed to explain neural models for text representation / indexing and ranking. In order to measure advances in this area, an evaluation framework is crucial. Humans constitute the primary target of explanations; the diversity of this target group makes it challenging to formulate an evaluation framework. The consumers of explanations range from lay users, through persons responsible for checking compliance with ethical / legal requirements, and domain experts who may not be familiar with the principles underlying IR / NLP models, to IR / NLP researchers and practitioners, whose primary objective may be failure analysis, or to gain insights that guide further development. Of course, this diversity is also true of the users of IR / NLP systems, and the original Cranfield paradigm for evaluating IR had several drawbacks. Some of these have been addressed, while some remain, but over the years, a generally accepted protocol for evaluating IR / NLP systems has emerged. Importantly, a number of large test collections have been created for many different tasks, and their role in advancing the field cannot be overemphasised.
@@
Even though similar, widely accepted protocols and test collections for ExIR are yet to emerge, two broad evaluation paradigms may be identified. The first involves anecdotal evaluation by humans. This seems unavoidable, since most research groups do not have the resources to conduct carefully designed, large-scale user studies. While this sort of evaluation may be enough to establish the plausibility of explanation methods, it is not rigorous. The second approach to evaluation utilises different quantitative notions of similarity. When explanations are in the form of simple surrogates that mimic complex models (see Section 3.4), evaluation is typically based on measuring the similarity between the behaviours of the explanation and the complex model, using standard metrics like rank correlation. In the ExIR literature, these metrics are generally termed fidelity measures. An explanation is considered faithful if it precisely mimics the underlying model, or accurately captures the reasoning underlying the model’s decision.
Likewise, feature attribution is generally evaluated by measuring the difference in the performance of a model when it is given the original inputs, and when it is given a simplified representation consisting of just the putatively important features. In this context, sufficiency and comprehensibility (deyoung-etal-2020-eraser) are two metrics that are widely used in the explainability literature. The sufficiency metric quantifies the extent to which the explanation component alone results in an accurate prediction. In contrast, comprehensibility measures the degree to which removing the explanation component causes a significant change in the model’s output. Intuitively, a high score on either metric indicates that the explanation component plays an important role in the model’s prediction process. Rationale-based approaches may be evaluated similarly. In cases where human annotated rationales are available, explanations may also be evaluated by measuring the overlap between the generated rationales and the ground truth. These metrics are often referred to as correctness measures.
Despite the existence of these metrics, evaluation was the focus of a long and animated discussion at the SIGIR 2025 workshop on Explainability in Information Retrieval (xir-workshop-sigir-25), and the lack of a definitive evaluation framework for ExIR was generally acknowledged. Thus, this is arguably the biggest issue facing the community, and needs to be resolved urgently.
@@ 6. Document Ranking
Pre LLM era Neural models for document ranking are generally categorised as representation-based or interaction-based. Like traditional retrieval models, representation-based models also encode queries and documents as vectors; however, in contrast to the sparse vectors used by traditional approaches, these vectors are dense. On the other hand, interaction-based models capture complex interactions across query-term–document-term pairs. This section provides an overview of efforts to explain both categories of neural ranking methods, but first, we provide a brief review of some foundational attempts to understand traditional IR models (Section 6.1). These early studies were all targeted towards global explanations. Next, in Section 6.2, we discuss the Axiomatic Framework (fang-sigir-04) — a cornerstone of a substantial body of work on explaining IR — and its recent applications. In Sections 6.36.5, we review other approaches to providing global explanations of Neural Ranking Models (NRMs). In Sections 6.66.9, we turn to studies devoted to local explanations. Tables 3 and 4 summarise the articles covered in these sections, according to whether they address global or local approaches. We have also prepared a Python package, ir_explain (ir-explain-resource) [^4] Available to use from https://github.com/souravsaha/ir_explain, that provides implementations of well-known post-hoc explanation strategies (including pointwise, pairwise, and listwise methods) within a single framework.
@@ 6.1. Understanding traditional models: the vector space model, and language modeling
Traditional IR systems typically represent documents using terms, i.e., single words (or word stems), and sometimes, additionally, groups of words that have special significance (these are often loosely referred to as phrases), along with their associated weights. The term weights (or feature weights) are calculated using either intuitive heuristics, or on the basis of various mathematical models. Traditional IR models may thus be regarded as explainable by design. Nonetheless, some early work (amit-thesis; kamps2005importance; smucker2005investigation) focused on developing a better understanding of these models. Since these studies try to provide general characterisations of specific models, in terms of the criteria introduced in Section 3, their explanations may be termed model-aware and global. We summarise these early efforts below.
amit-thesis carefully examined traditional term-weighting formulae used in the Vector Space Model, explained why they do not work as well as more effective statistical models like BM25, proposed ways to modify the formulae in order to address these flaws, and demonstrated that the modified formulae do indeed yield significant improvements in retrieval effectiveness. Besides conventional text collections, amit-thesis also considered noisy text collections, such as those generated by running Optical Character Recognition (OCR) on scanned document images, and explained why it is inadvisable to use cosine normalisation for such collections. kamps2005importance studied similar issues within the context of retrieving XML elements from a collection of XML documents, to explain the importance of length normalisation in XML retrieval. On a related note, smucker2005investigation compared two well-known smoothing techniques used in Language Modeling, and presented a different explanation for why Dirichlet smoothing outperforms Jelinek-Mercer smoothing, compared to the perspective provided by zhai2004study. We conclude this section by mentioning I-REX (i-rex), an interactive tool that is intended to facilitate further studies along these lines by providing a white-box view into traditional IR models. For example, given a retrieval model, and a query $Q$, the compare command in I-REX provides a pairwise explanation (Section 3.3) of the ranks of documents $D_{i},D_{j}$, by displaying the query terms present in $D_{i},D_{j}$ along with their term frequencies, collection frequencies, and overall contribution to the query-document scores.
@@ 6.2. The axiomatic framework: improving and explaining IR models
There are certain accepted principles (or heuristics) underlying the different term-weighting formulae that are commonly used, even though they are derived using different mathematical models. For example, it has been generally observed that documents having multiple query term matches are more likely to be relevant, and should therefore be ranked above documents that have a very strong match on only one term. In a series of studies (fang-sigir-04; semantic-term-matching-fang; fang-tao-zhai-tois-11), fang-sigir-04 consolidated these principles, and formalised them within an axiomatic framework, as a set of constraints, or axioms, that term-weighting schemes must satisfy, in order to be effective. This framework provides a natural basis for pairwise explanations (Section 3.3), i.e., for reasoning about why a particular document $D_{i}$ is ranked above $D_{j}$ by a ranking model for a given query $Q$. It has been used to explain why certain retrieval models work better than other models, as well as to formulate improved versions of existing retrieval models by better aligning them with IR axioms. (log-logistic-model; clinchant2009bridging). Later efforts have focused on refining the framework by incorporating additional axioms (proximity-measure-ir; semantic-term-matching-fang), as well as expanding it to cover pseudo-relevance feedback methods (clinchant2013theoretical). Table 2 lists some of the most popular axioms that have been mentioned in the literature. The page at https://www.eecis.udel.edu/~hfang/AX.html is an excellent resource on this topic.
@@
SemanticMatches STMC1 (semantic-term-matching-fang) favour a document containing terms that are more semantically related to query terms
@@
STMC2 (semantic-term-matching-fang) avoid over-favouring semantically similar terms
@@
STMC3 (semantic-term-matching-fang) favour a document containing more terms that are semantically related to different query terms
@@
TSFC (Pang2017ADI) a document having higher exact match should get higher score if their semantic similarity score matches
@@
Term Proximity TP (proximity-measure-ir) favour a document in which query terms appear in closer proximity
@@
Table 2. Popular axioms proposed in the past literature on axiomatic IR research (Section 6.2), adapted from (fang-tao-zhai-tois-11; camara-hauff-ecir).
@@
The axiomatic framework has been applied in recent times to analyse and understand neural IR models in terms of the common principles underlying traditional IR models (which are, in a sense, explainable by design). Some of these efforts (e.g., (rennings-diagnostic-dataset)) investigate the extent to which various IR models (neural, as well as traditional) satisfy axiomatic constraints. Others (e.g., (volske2021towards)) try to approximate the ranking generated by a retrieval model via an explainable surrogate, a la LIME (Section 4.2.1). In the rest of this section, we discuss (in chronological order) these attempts to study NRMs using the axiomatic framework.
@@
Pang2017ADI investigated the difference between the manually engineered features used in traditional models (BM25 and LMJM) and automatically learned features derived from two pre-BERT NRMs: ARC-I (arc-i), a representation based model, and MatchPyramid (matchpyramid), an interaction based model. Their experiments on the Robust04 and LETOR 4.0 (letor) datasets suggest that features automatically learned by NRMs are more robust compared to the word-based features used by BM25 and LMJM, in the sense that the performance of NRMs is less affected than that of traditional models when some features (according to the order of feature importance score) are removed. They also confirmed that interaction-based models appear to focus on query related words, while representation-based models seem more concerned with general topic related words, such as those generated by a state-of-the-art LDA based model. Interestingly, Pang2017ADI noted that ARC-I and MatchPyramid performed poorly on LETOR 4.0 and even worse on Robust04, compared to BM25 and LMJM. Their failure analysis suggests that these NRMs do not satisfy the TDC axiom (Table 2). In addition, due to limitations on memory and time, documents were sometimes truncated to some fixed length, e.g., 500 words. This truncation may reduce the number of query terms present in a document, resulting in a violation of the TFC1 axiom. It was also suggested that semantic matching may be introducing too much noise into the scoring function. A new axiomatic constraint TSFC (Table 2) was proposed, which states that if two documents are of equal length, and have equivalent semantic term matching signals, the document having a larger exact match should be ranked higher. Furthermore, the authors argued that the size of Robust04 is not suitable for training NRMs.
@@
Following the trend of creating diagnostic datasets for NLP and Computer Vision applications, rennings-diagnostic-dataset used selected axioms to construct a similar dataset for IR. Query-document-pair triplets that satisfy TFC1, TFC2, TDC (cf. Table 2) were sampled from the WikiPassageQA (wiki-passage-qa) corpus. Artificial documents were constructed to create triplets satisfying LNC2. This diagnostic dataset was used to evaluate the axiomatic efficacy (the fraction of axiomatic constraints that is satisfied by a retrieval model) of both traditional models (BM25 and LMDIR), and neural models like DRMM (drmm), aNMM (anmm), Duet (duet) and MatchPyramid (matchpyramid). The authors observed that statistical models outperform neural models in terms of axiomatic efficacy, but were significantly outperformed by DRMM and aNMM in terms of retrieval effectiveness. In later work, camara-hauff-ecir extended this idea by considering five additional axioms (LNC1, STMC1, STMC2, STMC3, TP). Once again, empirical results showed that DistilBERT (Sanh2019DistilBERTAD) significantly outperforms traditional query likelihood methods, but it does not perform as well in terms of axiomatic efficacy. It is particularly counter-intuitive that this observation applies to semantic matching axioms, e.g., STMC1, as well. These results suggest that existing axioms, including the semantic matching axioms, are not adequate for analysing NRMs based on BERT and its variants, and need to be formulated more carefully.
@@
volske2021towards provided post-hoc explanations of IR systems by constructing an explanation model to approximate the ranking generated by a retrieval model. Recall that each axiom induces relative orderings for certain document pairs; e.g., axiom $A_{i}$ may suggest that $D_{j}$ should be ranked above $D_{k}$. The explanation model (a random forest) may be regarded as a weighted aggregation of these orderings. In the words of the authors, “(a) the explanation model’s parameters reveal the degree to which different axiomatic constraints are important to the retrieval model under consideration, and (b) the fidelity with which the initial ranking can be reconstructed can point to blind spots in the axiom set, which can help to uncover new ranking properties yet to be formalized.” Twenty axioms were used in this study. The behaviour of three statistical models (BM25, TF-IDF, PL2) and five neural models (MP-COS (matchpyramid), DRMM (drmm), PACRR-DRMM (pacrr-drmm), BERT-3S (bert-3s), DAI-MAXP (dai-maxp)) were studied using the MSMARCO (msmarco) and Robust04 collections (please see Table 6 for a summary of information about some datasets commonly used in IR explainability studies). The authors observed that it is difficult to explain the relative ordering of two documents if they have similar retrieval scores, but a big difference in scores can be reliably explained in terms of the axioms. They also noted that explanation effectiveness is comparable for traditional and neural models (in this study, explanation efficacy is measured in terms of how accurately the explanation model can replicate a retrieval model’s ranking), but it was higher for the Robust04 collection, as compared to MSMARCO.
@@
The use of the axiomatic framework in these studies (rennings-diagnostic-dataset; camara-hauff-ecir; volske2021towards) draws on earlier work by hagen-cikm-16 that is not directly related to explainability. hagen-cikm-16 were likely the first to formulate relaxed versions of axioms that permitted them to be used for practical analysis,. Axiom TFC1 (cf. Table 2), for example, induces an ordering on a document pair based on the term-frequencies of a query term in the two documents, provided the two documents are of equal length. To be applicable in practice, this equality condition was replaced by an approximate equality constraint: axiom TFC1 was applied to a document pair if the relative difference in the lengths of the two documents was at most 10%. The work discussed above (rennings-diagnostic-dataset; camara-hauff-ecir; volske2021towards) uses such relaxed formulations of various axioms. It is worth noting that, even after the relaxations mentioned above, a number of the axioms (especially the later additions) involve pre-conditions that do not hold for most documents (or document pairs) retrieved by real models from benchmark collections. Thus, these axioms can rarely be applied in practice. Future research should focus on developing more representative axioms and rigorous test suites to better ground axiomatic principles in practical retrieval scenarios.